Assume the returns from holding small-company stocks are normally distributed. Also assume the average annual return for holding the small-company stocks for a period of time was 15.4 percent and the standard deviation of those stocks for the period was 33.3 percent. Use the NORMDIST function in Excel® to answer the following questions.


What is the approximate probability that your money will double in value in a single year? (Do not round intermediate calculations and enter your answer as a percent rounded to 3 decimal places, e.g., 32.161.)

Probability             %


What is the approximate probability that your money will triple in value in a single year? (Do not round intermediate calculations and enter your answer as a percent rounded to 8 decimal places, e.g., 32.16161616.)

Probability             %
Answer #1

Hello Sir/ Mam

Given that:

Mean = 15.40
Standard Deviation =                            33.30

Now,

(a)

100 - 15.4 P(X > 100) = 1-P(X < 100) = 1-P(:< -) = 1-P(Z < 33.3 P(X > 100) = 1 - P(Z < 2.540541) = 1 -0.9945 = 0.005499 = 0.5

(b)

P(X > 200) = 1-P(X < 200) = 1–P(25*- ) = 1–P(252 200 - 15.4 33.3 P(X > 200) = 1- P(Z < 5.543544) = 1-1=0 = 0%

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.